OpenStudy (anonymous):

Sam is building a new deck. His decking material is sold by the square foot. At a cost of $3.90 per square foot, how much will it cost Sam to build a deck that is 18 feet by 14 feet?

OpenStudy (anonymous):

as giving area=18*14=252 squrefoot so cost=3.9*252